√(2KP/R^3)^2 + √(KP/R^3)^2


1
Expert's answer
2021-05-13T18:08:24-0400
"\sqrt{\left(\frac{2KP}{R^3}\right)^2}+\sqrt{\left(\frac{KP}{R^3}\right)^2}\\=\frac{2KP}{R^3}+\frac{KP}{R^3}\\=\frac{3KP}{R^3}"